# A Stochastic Genetic Interacting Particle Method for Reaction-Diffusion-Advection Equations

### v1

@@ 3 Error Analysis of the SGIP method for RDA Equation
$|\langle w,-\nabla\cdot(\mathbf{v}w)\rangle|$ $\leq G_{1}\|\mathbf{v}\|_{L^{\infty}}\|w\|_{L^{2}}^{2}+\frac{D}{2}\|\Delta w\|_{L^{2}}^{2}$ (25)
where $G_{1}$ depends on the diffusion coefficient $D$ and the domain regularity.
For the diffusion term:
$D\langle w,\Delta w\rangle$ $=-D\|\nabla w\|_{L^{2}}^{2}\leq 0$ (26)
Thus:
$\frac{d}{dt}\|w\|_{L^{2}}^{2}$ $\leq 2G_{1}\|\mathbf{v}\|_{L^{\infty}}\|w\|_{L^{2}}^{2}$ (27)
By Gronwall’s inequality:
$\|w(\Delta t)\|_{L^{2}}^{2}$ $\leq e^{2G_{1}\|\mathbf{v}\|_{L^{\infty}}\Delta t}\|w(0)\|_{L^{2}}^{2}$ (28)
So for small $\Delta t$:
$\|S_{\mathcal{A}}(\Delta t)u_{0}-S_{\mathcal{A}}(\Delta t)v_{0}\|_{L^{2}}$ $\leq(1+L_{A}\Delta t)\|u_{0}-v_{0}\|_{L^{2}}$ (29)
where $L_{A}=2G_{1}\|\mathbf{v}\|_{L^{\infty}}$.
For the reaction equation $\partial_{t}u=r(u)$, consider two solutions $u(t)$ and $v(t)$ with difference w=uvw=u-v. Then:
$\frac{dw}{dt}$ $=r(u)-r(v)$ (30)
Using the Lipschitz condition $|r(u)-r(v)|\leq L_{r}|u-v|$ in Assumption 1:
$\left|\frac{d}{dt}\|w\|_{L^{2}}^{2}\right|$ $=2\left|\int w(r(u)-r(v))dx\right|\leq 2L_{r}\|w\|_{L^{2}}^{2}$ (31)
By Gronwall’s inequality:
$\|w(\Delta t)\|_{L^{2}}^{2}$ $\leq e^{2L_{r}\Delta t}\|w(0)\|_{L^{2}}^{2}$ (32)
So for small $\Delta t$:
$\|S_{\mathcal{R}}(\Delta t)u_{0}-S_{\mathcal{R}}(\Delta t)v_{0}\|_{L^{2}}$ $\leq(1+L_{R}\Delta t)\|u_{0}-v_{0}\|_{L^{2}}$ (33)
where $L_{R}=2L_{r}$.
Combining both estimates:
$\|S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{R}}(\Delta t)S_{\mathcal{A}}(\Delta t)u_{n}\|_{L^{2}}$
$\quad\leq(1+L_{R}\Delta t)\|S_{\mathcal{A}}(\Delta t)\tilde{u}_{n}-S_{\mathcal{A}}(\Delta t)u_{n}\|_{L^{2}}$
$\quad\leq(1+L_{R}\Delta t)(1+L_{A}\Delta t)\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$
$\quad\leq(1+C_{1}\Delta t)\|\tilde{u}_{n}-u_{n}\|_{L^{2}}$ (34)
where $C_{1}=L_{A}+L_{R}+O(\Delta t)$. ∎
Lemma 2 (Splitting error estimate).
Under Assumption 1, the Lie splitting error satisfies:
$\mathbb{E}\,\|u^{\mathrm{Lie}}_{n+1}-u_{n+1}\|_{L^{2}}\leq C_{2}\Delta t^{2}$ (35)
where $C_{2}$ depends on the commutator $\mathcal{A},\mathcal{R}]$ and the solution regularity.
Proof.
Using the Baker-Campbell-Hausdorff formula [2, 5, 13] for operator splitting, we have the formal expansion:
$S_{\mathcal{R}}(\Delta t)\circ S_{\mathcal{A}}(\Delta t)-S_{\mathcal{A}+\mathcal{R}}(\Delta t)=\tfrac{\Delta t^{2}}{2}[\mathcal{R},\mathcal{A}]+\mathcal{O}(\Delta t^{3})$ (36)
where $\mathcal{R},\mathcal{A}]u=r^{\prime}(u)\mathcal{A}u-\mathcal{A}(r(u))$ is the commutator expression for the reaction and advection-diffusion operators.
For the specific operators $\mathcal{R}u=r(u)$ and $\mathcal{A}u=\nabla\cdot(\mathbf{v}u)+D\Delta u$, we compute the commutator explicitly:
$[\mathcal{R},\mathcal{A}]u$ $=r^{\prime}(u)\mathcal{A}u-\mathcal{A}(r(u))$
$=r^{\prime}(u)[\nabla(\mathbf{v}u)+D\Delta u]-\left[\nabla\cdot(\mathbf{v}r(u))+D\Delta r(u)\right]$
$=r^{\prime}(u)\nabla(\mathbf{v}u)+Dr^{\prime}(u)\Delta u$
$\quad-\left[r(u)\nabla\mathbf{v}+r^{\prime}(u)\mathbf{v}\cdot\nabla u+Dr^{\prime}(u)\Delta u+Dr^{\prime\prime}(u)|\nabla u|^{2}\right]$
$=r^{\prime}(u)\left[\nabla(\mathbf{v}u)-\mathbf{v}\cdot\nabla u\right]-r(u)\nabla\mathbf{v}-Dr^{\prime\prime}(u)|\nabla u|^{2}$
$=r^{\prime}(u)u\nabla\mathbf{v}-r(u)\nabla\cdot\mathbf{v}-Dr^{\prime\prime}(u)|\nabla u|^{2}$ (37)
Under Assumption 1, using Hölder’s inequality and Sobolev embedding theorem, we can obtain:
$\|[\mathcal{R},\mathcal{A}]u\|_{L^{2}}$ $\leq\|r^{\prime}(u)u\nabla\mathbf{v}\|_{L^{2}}+\|r(u)\nabla\mathbf{v}\|_{L^{2}}+D\|r^{\prime\prime}(u)|\nabla u|^{2}\|_{L^{2}}$
$\leq\|r^{\prime}(u)\|_{L^{\infty}}\|u\|_{L^{\infty}}\|\nabla\mathbf{v}\|_{L^{\infty}}|\Omega|^{1/2}+\|r(u)\|_{L^{\infty}}\|\nabla\mathbf{v}\|_{L^{\infty}}|\Omega|^{1/2}+D\|r^{\prime\prime}(u)\|_{L^{\infty}}\|\nabla u\|_{L^{4}}^{2}$
$\leq G_{2}(\|\mathbf{v}\|_{W^{1,\infty}},\|r\|_{C^{2}},\|u\|_{H^{s}})$ (38)
with the constant $G_{2}$ depending on $|\Omega|$, $D$, and the Sobolev embedding constants.
The local error bound of order $\Delta t^{2}$ follows immediately from the Baker-Campbell-Hausdorff formula. ∎
Lemma 3 (Density estimation error).
Under Assumptions 1 and 2, the density estimation error satisfies:
@@
$\mathbb{E}\left[\|\widetilde{u}^{\mathrm{D}}_{n+1}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{2}}\right]\leq C_{3}(N\Delta x^{d})^{-1/2}+C_{4}\Delta x$ (39)
where the constants $C_{3},C_{4}$ depend on the domain size $L$, dimension $d$, diffusion coefficient $D$, the total mass at $t_{n}$ $\widetilde{m}_{n}=\int_{\Omega}\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}$, and the regularity parameters from Assumption 1.
Proof.
Let $\widetilde{M}_{n}=\int_{\Omega}\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}$ be the computed total mass at time $t_{n}$, which contains error from previous steps. The density estimate in bin $j$ after the advection-diffusion step is:
$\hat{u}_{j}=\frac{\widetilde{m}_{n}\cdot K_{j}}{(\Delta x)^{d}}=\frac{\widetilde{M}_{n}}{N(\Delta x)^{d}}\cdot K_{j}$ (40)
where $\widetilde{m}_{n}=\widetilde{M}_{n}/N$ is the particle mass at time $t_{n}$ and $K_{j}$ is the number of particles in bin $j$.
The density estimation error consists of two components: statistical sampling error, and spatial discretization error.
1. Statistical sampling error: Let $p_{j}^{\mathrm{true}}=\int_{\mathrm{bin}_{j}}\frac{S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})}{\widetilde{M}_{n}}d\mathbf{x}$ be the probability that a particle, starting from the initial distribution $\widetilde{u}_{n}$ and following the exact advection-diffusion dynamics, falls in bin $j$.
More precisely, consider particles $\mathbf{X}^{i}(t)$ satisfying the exact SDE:
$d\mathbf{X}^{i}(t)=\mathbf{v}(\mathbf{X}^{i}(t),t)dt+\sqrt{2D}d\mathbf{W}_{i}(t),\quad\text{for }t\in[t_{n},t_{n+1}]$ (41)
with initial distribution $\mathbf{X}^{i}(t_{n})\sim\widetilde{u}_{n}$. Then:
$p_{j}^{\mathrm{true}}=\mathbb{P}\left(\mathbf{X}^{i}(t_{n+1})\in\mathrm{bin}_{j}\mid\mathbf{X}^{i}(t_{n})\sim\widetilde{u}_{n}\right)$ (42)
The expected number of particles in bin $j$ is $\mathbb{E}[K_{j}]=Np_{j}^{\mathrm{true}}$, so the expected density is:
$\mathbb{E}[\hat{u}_{j}]=\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\cdot\mathbb{E}[K_{j}]=\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\cdot Np_{j}^{\mathrm{true}}=\frac{\widetilde{m}_{n}\cdot p_{j}^{\mathrm{true}}}{(\Delta x)^{d}}$ (43)
The variance of the density estimate is:
$\mathrm{Var}(\hat{u}_{j})=\left(\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\right)^{2}\cdot\mathrm{Var}(K_{j})\leq\left(\frac{\widetilde{m}_{n}}{N(\Delta x)^{d}}\right)^{2}\cdot Np_{j}^{\mathrm{true}}(1-p_{j}^{\mathrm{true}})$ (44)
Using $p_{j}^{\mathrm{true}}\leq\frac{\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}(\Delta x)^{d}}{\widetilde{m}_{n}}$, we get:
$\mathrm{Var}(\hat{u}_{j})\leq\frac{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}}{N(\Delta x)^{d}}$ (45)
Summing over all bins:
$\mathbb{E}\left[\sum_{j=1}^{M^{d}}\Delta x^{d}|\hat{u}_{j}-\mathbb{E}[\hat{u}_{j}]|^{2}\right]$ $=\sum_{j=1}^{M^{d}}\Delta x^{d}\cdot\mathrm{Var}(\hat{u}_{j})$
$\leq\sum_{j=1}^{M^{d}}\Delta x^{d}\cdot\frac{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}}{N(\Delta x)^{d}}$
$\leq\frac{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}(2L)^{d}}{N(\Delta x)^{d}}$ (46)
Taking square roots:
$\left(\mathbb{E}\left[\sum_{j=1}^{M^{d}}\Delta x^{d}|\hat{u}_{j}-\mathbb{E}[\hat{u}_{j}]|^{2}\right]\right)^{1/2}\leq\frac{C_{3}}{\sqrt{N}\Delta x^{d/2}}$ (47)
where $C_{3}=\sqrt{\widetilde{m}_{n}\|S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{\infty}}(2L)^{d}}$.
2. Spatial discretization error: Let $u_{j}^{*}=S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x}_{j})$ be the true density at the $j^{th}$-bin center $\mathbf{x}_{j}$. The bias due to spatial discretization is:
$|\mathbb{E}[\hat{u}_{j}]-u_{j}^{*}|=\left|\frac{\widetilde{m}_{n}\cdot p_{j}^{\mathrm{true}}}{(\Delta x)^{d}}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x}_{j})\right|$ (48)
Note that:
$\frac{\widetilde{m}_{n}\cdot p_{j}^{\mathrm{true}}}{(\Delta x)^{d}}=\frac{1}{(\Delta x)^{d}}\int_{\mathrm{bin}_{j}}S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}$ (49)
By Taylor expansion and the smoothness of $S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}$ under Assumption 1:
$\left|\frac{1}{(\Delta x)^{d}}\int_{\mathrm{bin}_{j}}S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})d\mathbf{x}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x}_{j})\right|\leq G_{3}\Delta x$ (50)
Thus, the $L^{2}$ spatial bias is:
$\left(\sum_{j=1}^{M^{d}}\Delta x^{d}|\mathbb{E}[\hat{u}_{j}]-u_{j}^{*}|^{2}\right)^{1/2}\leq C_{4}\Delta x$ (51)
where $C_{4}$ depends on the second derivatives of $S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}$ and the domain size.
Combining both error components Eqs.(47)(51):
@@
$\mathbb{E}\left[\|\widetilde{u}^{\mathrm{D}}_{n+1}-S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}\|_{L^{2}}\right]\leq C_{3}(N\Delta x^{d})^{-1/2}+C_{4}\Delta x$ (52)
This completes the proof. ∎
Remark 2 (On the particle motion discretization error).
The particle motion discretization error from the Euler-Maruyama scheme, which would contribute a $\Delta t$ term to the density estimation error, is intentionally omitted from Lemma 3. This error arises from the discrepancy between the numerical particle positions $\mathbf{X}_{i}^{*}$ (following Euler-Maruyama) and the exact positions $\mathbf{X}_{i}^{\text{true}}(t_{n+1})$ (following the exact SDE dynamics).
However, this error can be treated as an accumulation term and is absorbed into the propagation error estimate in Lemma 1. Specifically, the $\Delta t$ coefficient in the propagation error bound $(1+C_{1}\Delta t)$ implicitly incorporates the effects of particle motion discretization.
Lemma 4 (Reaction integration error).
@@
$\varepsilon_{\mathrm{react}}=\|\widetilde{u}^{\mathrm{R}}_{n+1}-S_{\mathcal{R}}(\Delta t)(\widetilde{u}^{\mathrm{D}}_{n+1})\|_{L^{2}}\leq\begin{cases}0,&\text{(exact integration)}\\ C_{5}\Delta t^{p},&\text{(numerical integration, order $p$)}\end{cases}$ (53)
where $C_{5}$ depends on the Lipschitz constant of $r(u)$ and the boundedness of higher derivatives under Assumption 1.
Proof.
Case 1: Exact integration For reactions with closed-form solutions (e.g., FKPP), the algorithm implements the exact solution operator:
$\widetilde{u}^{\mathrm{R}}_{n+1}=S_{\mathcal{R}}(\Delta t)(\widetilde{u}^{\mathrm{D}}_{n+1})\quad\Rightarrow\quad\varepsilon_{\mathrm{react}}=0$ (54)
Case 2: Numerical integration For general reactions, using a numerical method of order $p$ (e.g., implicit Euler: $p=1$, Crank-Nicolson: $p=2$) yields:

### v2

@@ 3 Error Analysis of the SGIP method for RDA Equation
(25) $\frac{1}{2}\frac{d}{dt}\|w_{1}\|^{2}$ $=\langle w_{1},-\nabla\cdot(\mathbf{v}w_{1})\rangle+D\langle w_{1},\Delta w_{1}\rangle\leq C\|\nabla\cdot\mathbf{v}\|_{\infty}\|w_{1}\|^{2}$
Using Assumption 1(b) and Gronwall’s inequality, we obtain:
(26) $\|w_{1}\|\leq e^{K_{A}\Delta t}\|w\|,$
where $K_{A}$ depends on $M_{1}$ and $L_{v}$ defined in Assumption 1(b).
Next, the operator $\mathcal{P}_{h}$ represents the $L^{2}$-orthogonal projection onto the finite volume space (piecewise constant functions). By the Cauchy-Schwarz inequality, for a mesh cell $B_{j}$ with volume $|B_{j}|$, we have:
(27) $\int_{B_{j}}|\mathcal{P}_{h}w_{1}|^{2}\,d\mathbf{x}=|B_{j}|\left|\frac{1}{|B_{j}|}\int_{B_{j}}w_{1}(\mathbf{x})\,d\mathbf{x}\right|^{2}\leq\int_{B_{j}}|w_{1}(\mathbf{x})|^{2}\,d\mathbf{x}.$
Summing over all cells $j$ yields the contraction property:
(28) $\|\mathcal{P}_{h}w_{1}\|^{2}=\sum_{j}\int_{B_{j}}|\mathcal{P}_{h}w_{1}|^{2}\,d\mathbf{x}\leq\sum_{j}\int_{B_{j}}|w_{1}|^{2}\,d\mathbf{x}=\|w_{1}\|^{2}.$
For the reaction step, let $w_{2}=\Phi_{h}(\tilde{u}_{n})-\Phi_{h}(\bar{u}_{n})$. By using the Lipschitz property of $r(u)$ with constant $L_{r}$ (Assumption 1(d)) and Gronwall’s inequality, we have:
(29) $\|w_{2}\|\leq e^{L_{r}\Delta t}\|\mathcal{P}_{h}w_{1}\|.$
Combining the estimates from all three steps, we obtain for sufficiently small $\Delta t$ that
$\|\Phi_{h}(\tilde{u}_{n})-\Phi_{h}(\bar{u}_{n})\|$ $\leq e^{(K_{A}+L_{r})\Delta t}\|\tilde{u}_{n}-\bar{u}_{n}\|$
(30) $\leq(1+C_{1}\Delta t)\|\tilde{u}_{n}-\bar{u}_{n}\|,$
where $C_{1}$ is a constant depending on the Lipschitz constants of the advection-diffusion and reaction operators. Thus, the proof is complete.
Lemma 3.7 (Consistency error estimate).
Under Assumption 1, the local consistency error between the ideal discrete scheme and the projected exact evolution satisfies:
(31) $\varepsilon_{\mathrm{consist}}=\|\Phi_{h}(\bar{u}_{n})-\bar{u}_{n+1}\|\leq C_{2}(\Delta t)^{2}+C_{3}\Delta t\Delta x,$
where $C_{2}$ and $C_{3}$ depend on the solution regularity and operator bounds.
Proof 3.8.
Recall $\Phi_{h}=S_{\mathcal{R}}\mathcal{P}_{h}S_{\mathcal{A}}$ and $\bar{u}_{n+1}=\mathcal{P}_{h}S_{\mathcal{R}+\mathcal{A}}u_{n}$. To facilitate the analysis, we introduce the intermediate reference state $\bar{u}_{n+1}^{*}=\mathcal{P}_{h}S_{\mathcal{R}+\mathcal{A}}(\Delta t)\bar{u}_{n}$, which represents the projected exact evolution starting from $\bar{u}_{n}$. We decompose the total consistency error into three components:
(32) $\varepsilon_{\mathrm{consist}}$ $\leq\|\Phi_{h}(\bar{u}_{n})-\mathcal{P}_{h}S_{\mathcal{R}}S_{\mathcal{A}}(\Delta t)\bar{u}_{n}\|+\|\mathcal{P}_{h}S_{\mathcal{R}}S_{\mathcal{A}}(\Delta t)\bar{u}_{n}-\bar{u}_{n+1}^{*}\|+\|\bar{u}_{n+1}^{*}-\bar{u}_{n+1}\|.$
(1) Time Splitting Error. The second term represents the classical Lie-Trotter splitting error. From Lemma 3.5, we have:
(33) $\|\mathcal{P}_{h}S_{\mathcal{R}}S_{\mathcal{A}}(\Delta t)\bar{u}_{n}-\mathcal{P}_{h}S_{\mathcal{R}+\mathcal{A}}(\Delta t)\bar{u}_{n}\|\leq\|S_{\mathcal{R}}S_{\mathcal{A}}(\Delta t)\bar{u}_{n}-S_{\mathcal{R}+\mathcal{A}}(\Delta t)\bar{u}_{n}\|$
By the Baker-Campbell-Hausdorff formula, this error is bounded by the commutator of the operators, which is defined as:
(34) $\mathcal{R},\mathcal{A}]u=r^{\prime}(u)\mathcal{A}u-\mathcal{A}(r(u))$
Expanding the term $\mathcal{A}(r(u))$ using the chain rule reveals that the non-vanishing commutativity arises primarily from spatial derivatives of the nonlinearity, introducing terms proportional to $r^{\prime\prime}(u)|\nabla u|^{2}$. Invoking Assumption 1(b) for velocity boundedness, Assumption 1(d) for the bound $|r^{\prime\prime}(u)|\leq M_{2}$, and the regularity of the solution, we obtain:
(35) $\|S_{\mathcal{R}}S_{\mathcal{A}}\bar{u}_{n}-S_{\mathcal{R}+\mathcal{A}}\bar{u}_{n}\|\leq\frac{1}{2}(\Delta t)^{2}\|[\mathcal{R},\mathcal{A}]\bar{u}_{n}\|\leq C_{\text{split}}(\Delta t)^{2}.$
Consequently, the splitting error is bounded by $C_{\text{split}}(\Delta t)^{2}$, where $C_{\text{split}}$ depends on $D$, the operator norms, and solution regularity.
(2) Spatial Consistency Error. The first term in Eq.(32) quantifies the error arising from the discretization of the reaction step on the grid. Specifically, it measures the discrepancy between applying the reaction operator on the projected density versus projecting the reacted density.
Let $v(\mathbf{x})=S_{\mathcal{A}}(\Delta t)\bar{u}_{n}(\mathbf{x})$ denote the density field after the exact advection-diffusion step but before projection. The spatial consistency error is exactly:
(36) $\|\Phi_{h}(\bar{u}_{n})-\mathcal{P}_{h}S_{\mathcal{R}}S_{\mathcal{A}}(\Delta t)\bar{u}_{n}\|$ $=\|S_{\mathcal{R}}(\Delta t)(\mathcal{P}_{h}v)-\mathcal{P}_{h}(S_{\mathcal{R}}(\Delta t)v)\|.$
Since the reaction function $r(u)$ is twice continuously differentiable (Assumption 1(d)), we can expand the exact reaction evolution $S_{\mathcal{R}}(\Delta t)u$ via Taylor series with an integral remainder. For any state $w$, we have:
(37) $S_{\mathcal{R}}(\Delta t)w=w+\Delta tr(w)+\mathcal{E}_{t}(w),$
where $\|\mathcal{E}_{t}(w)\|\leq K_{r}(\Delta t)^{2}$ and the constant $K_{r}$ depends on the bounds of $r$ and its derivatives.
Applying this expansion to both terms in the error expression:
(38) $S_{\mathcal{R}}(\Delta t)(\mathcal{P}_{h}v)$ $=\mathcal{P}_{h}v+\Delta tr(\mathcal{P}_{h}v)+\mathcal{E}_{t}(\mathcal{P}_{h}v),$
(39) $\mathcal{P}_{h}(S_{\mathcal{R}}(\Delta t)v)$ $=\mathcal{P}_{h}\left(v+\Delta tr(v)+\mathcal{E}_{t}(v)\right)=\mathcal{P}_{h}v+\Delta t\mathcal{P}_{h}r(v)+\mathcal{P}_{h}\mathcal{E}_{t}(v).$
Subtracting the two equations and using the triangle inequality yields:
(40) $\|\!S_{\mathcal{R}}(\Delta t)(\mathcal{P}_{h}v)\!-\!\mathcal{P}_{h}(S_{\mathcal{R}}(\Delta t)v)\|\!\leq\!\Delta t\|r(\mathcal{P}_{h}v)\!-\!\mathcal{P}_{h}r(v)\|\!+\!\|\mathcal{E}_{t}(\mathcal{P}_{h}v)\|\!+\!\|\mathcal{P}_{h}\mathcal{E}_{t}(v)\|.$
The first term $\|r(\mathcal{P}_{h}v)-\mathcal{P}_{h}r(v)\|$ represents the commutator between the nonlinear reaction function and the projection operator. Using the Lipschitz property of $r(u)$ with constant $L_{r}$ and the approximation properties of the projection $\mathcal{P}_{h}$, this term is bounded by
(41) $\|r(\mathcal{P}_{h}v)-\mathcal{P}_{h}r(v)\|\leq C_{P}\Delta x\|\nabla r(v)\|\leq C_{P}L_{r}\Delta x\|\nabla v\|.$
Crucially, $v=S_{\mathcal{A}}(\Delta t)\bar{u}_{n}$ is the result of the advection-diffusion operator acting on the piecewise constant function $\bar{u}_{n}$. Due to the smoothing property of the parabolic operator for $t>0$, $v$ belongs to $H^{1}(\Omega)$. Moreover, given the smoothness of the exact solution (Assumption 1), $\|\nabla v\|$ remains uniformly bounded.
For the remainder terms, since the projection operator is non-expansive in $L^{2}$, we have:
(42) $\|\mathcal{E}_{t}(\mathcal{P}_{h}v)\|+\|\mathcal{P}_{h}\mathcal{E}_{t}(v)\|\leq 2K_{r}(\Delta t)^{2}.$
Combining these bounds, we obtain:
(43) $\|\Phi_{h}(\bar{u}_{n})-\mathcal{P}_{h}S_{\mathcal{R}}S_{\mathcal{A}}(\Delta t)\bar{u}_{n}\|\leq C_{\text{spatial}}\Delta t\Delta x+2K_{r}(\Delta t)^{2},$
where $C_{\text{spatial}}=C_{P}L_{r}\|\nabla u\|_{L^{\infty}(0,T;L^{2})}$.
(3) Projection Stability Error. The last term accounts for the evolution of the initial projection error over one time step. Using the first-order Taylor expansion of the exact evolution operator $S_{\mathcal{R}+\mathcal{A}}(\Delta t)=I+\Delta t(\mathcal{R+A})+O(\Delta t^{2})$, we have:
$\|\bar{u}_{n+1}^{*}-\bar{u}_{n+1}\|$ $=\|\mathcal{P}_{h}((I+\Delta t\mathcal{L})\bar{u}_{n}-(I+\Delta t\mathcal{L})u_{n})+O(\Delta t^{2})\|$
(44) $\leq\|\mathcal{P}_{h}(\bar{u}_{n}-u_{n})\|+\Delta t\|\mathcal{P}_{h}\mathcal{L}(\bar{u}_{n}-u_{n})\|+C_{\text{rem}}(\Delta t)^{2}.$
where we define $\mathcal{L}=\mathcal{R+A}$. Crucially, since $\bar{u}_{n}=\mathcal{P}_{h}u_{n}$ is the projection of $u_{n}$, we have $\mathcal{P}_{h}(\bar{u}_{n}-u_{n})=\mathcal{P}_{h}(\mathcal{P}_{h}u_{n}-u_{n})=0$. The remaining term is bounded by the Lipschitz continuity of the generator $\mathcal{L}$ and the projection error bound $\|\bar{u}_{n}-u_{n}\|\leq C\Delta x$:
(45) $\Delta t\|\mathcal{P}_{h}\mathcal{L}(\bar{u}_{n}-u_{n})\|\leq C_{\mathcal{L}}\Delta t\|\bar{u}_{n}-u_{n}\|\leq C_{\text{stab}}\Delta t\Delta x.$
Combining the estimates from Eqs.(35)(43)(3.8), we obtain:
(46) $\varepsilon_{\mathrm{consist}}$ $\leq(C_{\text{split}}+2K_{r}+C_{\text{rem}})(\Delta t)^{2}+(C_{\text{spatial}}+C_{\text{stab}})\Delta t\Delta x.$
By defining constants $C_{2}=C_{\text{split}}+2K_{r}+C_{\text{rem}}$ and $C_{3}=C_{\text{spatial}}+C_{\text{stab}}$, we obtain the stated bound in Eq.(31).
Lemma 3.9 (Statistical density estimation error).
Under Assumption 1, the density estimation step is unbiased with respect to the ideal discrete state, and its mean squared error satisfies:
@@
(47) $\mathbb{E}\left[\|\varepsilon_{\mathrm{dens}}\|^{2}\right]=\mathbb{E}\!\left[\big\|S_{\mathcal{R}}(\Delta t)(\widetilde{u}^{\mathrm{D}}_{n+1})-S_{\mathcal{R}}(\Delta t)(\mathcal{P}_{h}S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n})\big\|^{2}\right]\leq\frac{C_{4}}{N(\Delta x)^{d}},$
where $C_{4}$ depends on the domain size and the uniform bound of the solution.
Proof 3.10.
Let $v(\mathbf{x})=S_{\mathcal{A}}(\Delta t)\widetilde{u}_{n}(\mathbf{x})$ denote the advected-diffused density field. The reference state is its projection $\bar{v}_{j}=\frac{1}{(\Delta x)^{d}}\int_{\text{bin}_{j}}v(\mathbf{x})d\mathbf{x}$. The numerical density in bin $j$, $\hat{u}_{j}$, is constructed from $N$ independent particle positions $\{\mathbf{X}_{i}\}_{i=1}^{N}$ sampled from the distribution $v(\mathbf{x})/M$, where $M$ is the total mass.
Let $I_{j}$ be the number of particles in bin $j$. The variable $I_{j}$ follows a binomial distribution $I_{j}\sim\text{Binomial}(N,p_{j})$, where the probability $p_{j}$ is given by:
(48) $p_{j}=\int_{\text{bin}_{j}}\frac{v(\mathbf{x})}{M}d\mathbf{x}=\frac{(\Delta x)^{d}\bar{v}_{j}}{M}.$
The estimator is defined as $\hat{u}_{j}=\frac{MI_{j}}{N(\Delta x)^{d}}$. Taking the expectation:
(49) $\mathbb{E}[\hat{u}_{j}]=\frac{M}{N(\Delta x)^{d}}\mathbb{E}[I_{j}]=\frac{M}{N(\Delta x)^{d}}(Np_{j})=\frac{M}{(\Delta x)^{d}}\frac{(\Delta x)^{d}\bar{v}_{j}}{M}=\bar{v}_{j}.$
Thus, $\mathbb{E}[\hat{u}_{j}]-\bar{v}_{j}=0$, confirming the estimator is unbiased with respect to the projected state.
The Mean Squared Error is the sum of variances in each bin:
(50) $\mathbb{E}[\|\widetilde{u}^{\mathrm{D}}_{n+1}-\bar{v}\|^{2}]=\sum_{j\in\mathcal{G}}(\Delta x)^{d}\mathrm{Var}(\hat{u}_{j}).$
Using the variance of the binomial distribution $\mathrm{Var}(I_{j})=Np_{j}(1-p_{j})$:
(51) $\mathrm{Var}(\hat{u}_{j})=\frac{M^{2}}{N^{2}(\Delta x)^{2d}}Np_{j}(1-p_{j})\leq\frac{M^{2}}{N(\Delta x)^{2d}}p_{j}.$
Substituting $p_{j}=\frac{(\Delta x)^{d}\bar{v}_{j}}{M}$ into the sum, we have:
@@
(52) $\mathbb{E}[\|\widetilde{u}^{\mathrm{D}}_{n+1}-\bar{v}\|^{2}]\leq\sum_{j\in\mathcal{G}}(\Delta x)^{d}\left(\frac{M^{2}}{N(\Delta x)^{2d}}\frac{(\Delta x)^{d}\bar{v}_{j}}{M}\right)\leq\frac{M^{2}}{N(\Delta x)^{d}}.$
Since the reaction term satisfies Assumption 1(d) with Lipschitz constant $L_{r}$, the associated solution operator $S_{\mathcal{R}}(\Delta t)$ is Lipschitz continuous on $L^{2}(\Omega)$ with constant $L_{\mathrm{R}}=e^{L_{r}\Delta t}\leq e^{L_{r}T}$. Therefore,
$\mathbb{E}\!\left[\varepsilon_{\mathrm{dens}}^{2}\right]\leq L_{\mathrm{R}}^{2}\,\mathbb{E}\!\left[\big\|\widetilde{u}^{\mathrm{D}}_{n+1}-\mathcal{P}_{h}v\big\|^{2}\right]\leq\frac{C_{4}}{N(\Delta x)^{d}},$
where $C_{4}=e^{2L_{r}T}(C_{M})^{2}$ depends on the final time $T$ and the mass bound $C_{M}$ established in Lemma 3.3.
Lemma 3.11 (Reaction integration error).
@@
(53) $\varepsilon_{\mathrm{react}}=\|\tilde{u}^{\mathrm{R}}_{n+1}-S_{\mathcal{R}}(\Delta t)(\tilde{u}^{\mathrm{D}}_{n+1})\|\leq C_{5}(\Delta t)^{p+1},$
where $p$ is the order of the time integration scheme ($p\to\infty$ for exact integration).
Proof 3.12.
Let $\Psi_{\Delta t}(\cdot)$ denote the numerical integration operator over one time step $\Delta t$, and $S_{\mathcal{R}}(\Delta t)$ be the exact reaction operator. The local truncation error for a $p$-th order method applied to the ODE $\dot{u}=r(u)$ is given pointwise by:
(54) $|\Psi_{\Delta t}(u(\mathbf{x}))-S_{\mathcal{R}}(\Delta t)u(\mathbf{x})|\leq C_{r}(\Delta t)^{p+1},$
where $C_{r}$ depends on the $(p+1)$-th derivative of the solution, which is bounded due to the smoothness of $r(u)$ and the boundedness of $u$ (Assumption 1 (c)(d)).
The $L^{2}$-norm of the error is: